A moon orbits a planet at a distance of 1,010,000 km the moons orbital period about the planet is 15.9 days. What is the mass of the planet?
1
Expert's answer
2017-02-01T14:00:15-0500
T^2=4Pi^2a^3/(GM), M=4Pi^2a^3/(GT^2)=4×3.14^2×1.01^3×10^27/(6.67×10^(-11)×[15.9×24×3600]^2)=3.22×10^26 kg